OpenStudy (anonymous):

How do I make an equation for this parabola knowing: Vertex: (3,2) y-intercept: -16 x-intercepts: 2 and 4

OpenStudy (sriram):

(x-h)^2=-4a(y-k) form h=3 k=2 put the points 4,0 2,0 & 0,-16 to get the value of a

OpenStudy (anonymous):

if your x-intercepts: 2 and 4 then (x-2)(x-4)=0 that is x^2 -6x+8=0, assuming y-intercept: +16 ,then multiplying by 2 to get 2x^2 -12x+16=0 Equation of parabola vertex v=-b/2a=-(-12)/2(2)=3 and f(3)=3^2 -12(3)+16=-2 therefore vertex(3,-2)